OpenStudy (anonymous):

1- 1/2 +1/4 + 1/5 +1/6 ?? plz help

OpenStudy (anonymous):

Just break it into smaller pieces so it's simpler. 1- 1/2 = 2/2 - 1/2 = 1/2 so now it's 1/2 + 1/4 + 1/5 + 1/6 take the next two 1/2 + 1/4 = 2/4 + 1/4 = 3/4 3/4 + 1/5 + 1/6 3/4 + 1/5 = 15/20 + 4/20 = 19/20 19/20 + 1/6 = 57/60 + 10/60 = 67/60